An old person cannot see objects closer than 200 cm from his eye. Determine the lens he should wear to view objects at 25 cm from his eye.
5 Leel Ashokarathne
If the rays converging at a point not less than 150 cm behind the eye lens can be focused by the eye, find the visual range with spectacles.
1
Expert's answer
2020-11-03T10:32:26-0500

The objects 200 cm away must be "shifted" by the lens to the distance of 25 cm from the eye. Thus, apply thin lens equation:


"f=\frac{oi}{o+i}=\frac{(-2)\cdot0.25}{-2+0.25}=0.286\text{ m}.\\\space\\
P=\frac{1}f=3.5D."

The person is far-sighted, converging lenses are needed.